The sum of two numbers is 28. Their difference is 4. What are the two numbers?
thanks so much for the help!

Answer by CharStar(110)

x= 1st number
x-4 = 2nd number
x + x-4 = 28
Subtract 4 from both sides and add the x
2x=32
x=16
Check you answer
x=16
x-4 = 12
16+12 = 28